Dear Richard dear all,

please find below a short summary of the activities of Sunday & Monday on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: P. Stefanini, E. Pinna, A. Riccardi, R. Briguglio, S. Esposito, L. Busoni, M. Xompero
ADS: P. Lazzarini

Total # of person: 8

W unit

No activity

AO control SW

1) Coding of the ASM calibration and test GUI for the telescope.

LBT672a

1) Final check of the alignment of the retroreflector structure. The internal alignment is inside specifications 1.4 mm decenter between top and bottom rings.
2) verification of the power on procedure and temperature setting of the box.
3) the secondary mirror is set in flatten position since 24h and is still set without any need for manual adjustement.

Logistic

Due to the snow storm on the mountain A. Riccardi has been not able to take is flight back to Italy. The new departure date of A: Riccardi is 17th of March

Next update tomorrow

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback